Diffusion Morning Report 02/18/09

Ted Berg tberg at stanford.edu
Wed Feb 18 06:42:54 PST 2009


Good Morning All,
    Looking OK. Busy this morning. Poly and TEOS 2 down for pull and 
clean. SiGe is in temp cal Ebeam still tweeking beam. No other issues at 
this time.

  	  	  	  	  	 
  	Date: 	2/18/09 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	  	  	  	 
Tylan 1 	up 	  	  	  	 
Tylan 2 	up 	  	  	  	 
Tylan 3 	up 	  	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	up 	  	  	  	 
Tylan 6-POCl3 	up 	  	  	  	 
Tylan 7-TEOS2 	down 	pull and clean 	0/30 	  	 
Tylan 8-SiGe 	yellow 	temp cal 	1.12/30 	  	 
Tylan 9-Poly 	down 	pull and clean 	0/30 	  	 
Tylan 10-Nitride 	up 	  	.27/30 	  	 
Tylan 11-LTO 	up 	  	25.63/30 	  	 
Tylan 12-FGA 	up 	  	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	up 	  	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	up 	  	  	  	 
Thermco Poly2 	up 	  	  	  	 
Thermco Nitride1 	up 	  	  	  	 
Thermco Nitride2 	future 	  	  	  	 
TEL Spa 	up 	  	  	  	 
Burn Box 	up 	  	   	  	 
Ebeam 	down 	tip changed looking for beam 	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	up 	  	  	  	 
SiH4 for furnaces 	  	B on line 60 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20090218/853d6293/attachment.html>


More information about the diffusion mailing list